OpenStudy (anonymous):

Divide. (4m^9 - 8m^8 - 14m^6) / 2m^2

OpenStudy (anonymous):

(4m^9 - 8m^8 - 14m^6) / 2m^2 = (2m^7 - 4m^6 - 7m^3)

OpenStudy (anonymous):

\[\frac{(4m^9 - 8m^8 - 14m^6)}{2m^2} = \frac{4m^9}{2m^2} - \frac{8m^8}{2m^2} - \frac{14m^6}{2m^2}\]

OpenStudy (anonymous):

\[2m^7 - 4m^6 - 7m^4\]

OpenStudy (anonymous):

Yepper, I made a miscalculation on the end (2m^7 - 4m^6 - 7m^4)

OpenStudy (earthcitizen):

\[m^4 (2 m^3-4 m^2-7)\]